Ask
your own question, for FREE!
Mathematics
18 Online
A shelf has 20 bags of potatoes. Two of the bags have potatoes starting to go bad. What is the expected probability of the complement of the event "bad bags of potatoes"? 0.1 0.2 0.8 0.9
Still Need Help?
Join the QuestionCove community and study together with friends!
0.1
.2? That is how many fell out of the bag.
Wait nvm
That are bad not how many fell out.
There are 2 bad bags out of 20, so the probability of bad bags is 2/10 = 1/10 = 0.1 But we want the complement, which is the probability of bags which are not bad, and that's 1 - 0.1 = d] 0.9 as discussed above.
Still Need Help?
Join the QuestionCove community and study together with friends!
0.1
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
stxandrew1214:
I lwk like a junior need help to talk with her and get to know her but I think sh
thatonenerd:
Use Hvap of water = 40.7 kJ/mol, c of water = 4.184 J/g.u00b0C, and molar mass of HO = 18.
letrasbonitasz:
Typography plays an important role in digital communication. Different font styles can make social media profiles, usernames, messages, and creative designs
Bounty:
tested love: im so over love it acquires patience i don't have trials an tribulations yet my souls the only one taken i give 90 you give 10 you say that you
SaltySuga:
is it possible to tell how deep a cut is by an open wound? like would i be able t
AndreW2002:
I'm trying to change my profile picture but it keeps saying "Either a error occurred, or you tried to upload a gif" but it's not a gif that I'm using, it's
15 hours ago
19 Replies
1 Medal
20 hours ago
5 Replies
0 Medals
2 days ago
3 Replies
1 Medal
19 hours ago
7 Replies
2 Medals
3 days ago
5 Replies
1 Medal
4 days ago
3 Replies
0 Medals